OpenStudy (anonymous):

Which of the following is a polynomial with roots -2, -3i, and 3i ?

OpenStudy (anonymous):

x3 + 4x2 + 9x + 24 x3 - 4x2 + 9x - 24 x3 + 2x2 + 9x + 18 x3 - 2x2 + 9x - 18

OpenStudy (anonymous):

It would be \[Answer=(x+2)(x+3i)(x-3i)\]

OpenStudy (anonymous):

how would you solve that?

OpenStudy (anonymous):

Multiply the first term by the second term. Then take the product and multiply that with the third and final term.

OpenStudy (anonymous):

Use the law of distribution.\[a*(b+c)=(a*b)+(a*c)\]

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

Here's a guide for using the FOIL method: http://www.algebrahelp.com/lessons/simplifying/foilmethod/

OpenStudy (anonymous):

i know how to foil

OpenStudy (anonymous):

Good! Use the foil method on the first two terms.

OpenStudy (anonymous):

Show me what you get.

OpenStudy (anonymous):

x^2... im stuck on the x*3i

OpenStudy (anonymous):

\[(x+2)(x+3i)=x^{2}+3x i+2x+6i\]

OpenStudy (anonymous):

Then multiply that by the last term.

OpenStudy (anonymous):

oh.. im dumb

OpenStudy (anonymous):

Not at all! We are all here to learn. That's not dumb.

OpenStudy (anonymous):

so what do i do form there

OpenStudy (anonymous):

\[(x^{2}+3x i+2x+6i)(x-3i)=(x^{2}+3x i+2x+6i)(x)-(x^{2}+3x i+2x+6i)(3i)\]If you look at my generic example of the law of distribution above, 'a' would be that long x squared number.

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

these problems suck.. i dont get it

OpenStudy (anonymous):

Can you solve (x2+3xi+2x+6i)(x)?

OpenStudy (anonymous):

I mean simplify, not solve technically.

OpenStudy (anonymous):

3x+4xi+3x+6xi

OpenStudy (anonymous):

What's x^2 times x?

OpenStudy (anonymous):

x^3

OpenStudy (anonymous):

Good. That's the first term. The next term would be 3xi times x.

OpenStudy (anonymous):

3i^2

OpenStudy (anonymous):

Remember that x and i are different variables representing different numbers. So it would be (3)(x)(i)(x)=3x^2i

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Next will be (2x)(x) + (6i)(x)

OpenStudy (anonymous):

3x+6xi

OpenStudy (anonymous):

(2x)(x)=\[2x^{2}\]

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

Okay. So what's \[(x^{2}+3x i +2x+6i)(x)=?\]

OpenStudy (anonymous):

x^3+3x2i+2x^2+6xi?

OpenStudy (anonymous):

Very good. Now do \[(x^{2}+3x i+2x+6i)(3i)\]

OpenStudy (anonymous):

3xi^2+6xi^2+6xi+18i^2

OpenStudy (anonymous):

Very close. Look at the 3xi times 3i again.

OpenStudy (anonymous):

im not sure on that one

OpenStudy (anonymous):

You added the three's instead of multiplying them

OpenStudy (anonymous):

oh 9xi

OpenStudy (anonymous):

\[9x i^{2}\]

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

It's okay. I've actually got to get ready for work tmrw. I'm sure someone here will help you finish. @satellite73 @TuringTest @ganeshie8

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

\[\large (x+2)(x+3i)(x-3i)\]

ganeshie8 (ganeshie8):

may be first multiply the last two factors

ganeshie8 (ganeshie8):

\[\large (x+2)\color{red}{(x+3i)(x-3i)}\]

ganeshie8 (ganeshie8):

recall the identity : \((a+b)(a-b)=a^2-b^2\)

ganeshie8 (ganeshie8):

\[\large (x+2)\color{red}{(x^2 - (3i)^2)}\]

OpenStudy (anonymous):

umm im kinda lost